Block A weighs 1.4 N and block B weighs 4.2 N. The coefficient of kinetic friction between all surfaces is 0.3.

Find the magnitude of force F as shown in Fig. 1 necessary to drag block B to the
left at an acceleration of 2 m/s2 if A and B

F А 30° B Fig. 1

are connected by a light flexible cord passing around a fixed frictionless pulley.
Answer #1

F Efu=0 EF= 30 appy 2 I watwa) 2 wt Fsin 30° - Wat we to > N = 1.444.2 - 0.5F = (s.6-0. Sf) MN-u (6.6-0.5 sp) fA=ff Efn - mBa